# Memory Is Not the Bottleneck: Cost-Efficient Continual Learning via Weight Space Consolidation

### v3

@@ A.4 Experimental Results and Analysis
Table 5: Average class-IL accuracy (%) on CIFAR-100 with varying exemplar-memory sizes. We report experimental results with varying memory sizes, ranging from 20 exemplars per class (a common setting in class-IL) to 500 exemplars per class (storing the entire previous dataset in CIFAR-100. Bold highlights the best non-expansion method.
@@
The results of the CIFAR-100 class-IL experiment are reported in Table 5. Here, we validate that our method is indeed the strongest among non-expansion methods, and even surpasses expansion-based methods under sufficient memory (see Figure 9(a)). Specifically, we see that in cases where the exemplar memory size is larger than $16\%$ of the full data, our weight space consolidation method outperforms all non-expansion methods, and begins to match the costly expansion-based methods when the exemplar memory ratio exceeds $40\%$. The true strength of our method lies in its training cost (see Figure 9(b)), where our method’s train time is at par with the cheapest baseline (Replay), while taking one-fifth the time of the expansion-based FOSTER [61] and non-expansion-based BiC [67].
@@
Table 6: Average class-IL accuracy (%) on ImageNet-100 with varying exemplar-memory sizes. Bold highlights the best non-expansion method.
@@
The experimental results in the ImageNet-100 benchmark display a similar pattern. In Figure 10(a), we observe a gradual increase in average task accuracy as the exemplar memory size increases, which eventually saturates as it enters the sufficient memory regime. The cost of training increases proportionally to the memory size, where methods like BiC or DER require substantially larger training time. On the other hand, our method displays high accuracy while using roughly one-third   half of the training time (see Figure 10(b)). For detailed results, please refer to Table 6.
@@
Table 7 reports the LLM continual instruction tuning benchmark results in TRACE. Similar to the results in the Class-IL benchmarks (e.g., CIFAR-100, ImageNet-100), we observe a similar pattern where the accuracy grows in accordance with the increase in memory size (see Figure 3). An interesting observation is that offline merging (i.e., model merging methods that merge post-training) tends to underperform, displaying lower scores while requiring the VRAM cost of storing multiple models. This aligns with the empirical results of Dziadzio et al. [15], which showed that offline merging methods face difficulties in accumulating the multi-task knowledge. Compared to this, our method In addition, note that for TRACE, we use a different method to measure the train cost, which is the relative VRAM usage. The relative VRAM usage is a scaled version of the training time, which measures the computation cost of the training based on the VRAM usage. This scaling is required to distinguish from models that take a similar time to train, but use different numbers of GPUs. For instance, our method and MagMax require similar time to train, but MagMax requires multiple ($t$) task vectors to be loaded in the VRAM, and this is reflected in the relative VRAM usage measure. Considering this, we visualize the results in Figure 4.
@@
Table 7: General language understanding and reasoning scores on TRACE with varying exemplar-memory sizes.
@@
Method # Mem. Size # of $\theta$ stored Runtime (min.) C-STANCE FOMC MeetingBank Py150 ScienceQA $\text{NumGLUE}_{\text{cm}}$ $\text{NumGLUE}_{\text{ds}}$ 20Minuten Avg. Acc. Acc. ROUGE-L Similarity Acc. Acc. Acc. SARI Fine-tune 137’5 0.42 0.25 0.33 0.49 0.21 0.28 0.51 0.37 0.35 Replay 0.2 314’7 0.43 0.01 0.35 0.50 0.29 0.32 0.38 0.37 0.33 0.4 351’2 0.49 0.21 0.38 0.50 0.47 0.35 0.41 0.39 0.40 0.6 394’7 0.47 0.13 0.38 0.55 0.18 0.40 0.56 0.37 0.38 0.8 449’4 0.49 0.13 0.37 0.57 0.21 0.38 0.56 0.37 0.38 SLERP 0.2 1 375’9 0.42 0.45 0.31 0.54 0.36 0.29 0.42 0.40 0.39 0.4 1 422’4 0.45 0.42 0.26 0.53 0.33 0.29 0.48 0.43 0.40 0.6 1 472’9 0.42 0.56 0.38 0.57 0.30 0.34 0.54 0.38 0.43 0.8 1 526’4 0.45 0.18 0.37 0.58 0.39 0.37 0.53 0.38 0.41 Avg. (Online) 0.2 1 374’5 0.44 0.46 0.33 0.52 0.39 0.31 0.37 0.37 0.40 0.4 1 421’7 0.42 0.41 0.35 0.55 0.43 0.28 0.43 0.44 0.41 0.6 1 471’4 0.45 0.50 0.41 0.55 0.35 0.34 0.52 0.39 0.44 0.8 1 524’5 0.45 0.49 0.35 0.52 0.16 0.35 0.47 0.36 0.39 Avg. (Offline) 0.2 $t$ 377’1 0.48 0.00 0.37 0.58 0.29 0.34 0.51 0.37 0.36 0.4 $t$ 430’6 0.45 0.03 0.38 0.58 0.33 0.40 0.48 0.37 0.38 0.6 $t$ 479’3 0.44 0.56 0.38 0.59 0.36 0.38 0.53 0.36 0.45 0.8 $t$ 525’4 0.50 0.21 0.38 0.59 0.35 0.38 0.49 0.37 0.41 Task-Arith. 0.2 $t$ 374’6 0.51 0.02 0.39 0.56 0.29 0.40 0.52 0.38 0.38 0.4 $t$ 427’6 0.49 0.01 0.39 0.59 0.32 0.40 0.53 0.39 0.39 0.6 $t$ 474’9 0.50 0.47 0.38 0.60 0.32 0.42 0.52 0.40 0.45 0.8 $t$ 523’1 0.49 0.28 0.39 0.57 0.40 0.44 0.48 0.38 0.43 MagMax 0.2 $t$ 373’3 0.49 0.44 0.25 0.26 0.37 0.31 0.48 0.39 0.37 0.4 $t$ 424’8 0.46 0.41 0.39 0.26 0.35 0.43 0.43 0.40 0.39 0.6 $t$ 469’9 0.59 0.45 0.37 0.32 0.43 0.42 0.45 0.40 0.42 0.8 $t$ 525’6 0.32 0.54 0.23 0.25 0.32 0.29 0.51 0.39 0.36 Ours 0.2 1 370’9 0.39 0.13 0.36 0.51 0.32 0.40 0.56 0.39 0.38 0.4 1 422’6 0.41 0.48 0.38 0.55 0.45 0.41 0.30 0.39 0.42 0.6 1 465’8 0.44 0.57 0.35 0.52 0.47 0.38 0.55 0.39 0.46 0.8 1 522’9 0.46 0.47 0.37 0.55 0.46 0.36 0.53 0.40 0.45
@@
Lastly, we show that under sufficient exemplar memory, convergence becomes difficult, similar to a multi-task learning setting. The results are illustrated in Figure 2(b). Specifically, we analyze the training loss of the model under two cases. (1) Continual Learning: train the model across 20 tasks sequentially, with sufficient exemplar memory. (2) Joint Training: train individual models for each task using all known task data. For this experiment, we used the CIFAR-100 dataset divided into 20 subtasks. In Figure 2(b), we can see that under sufficient exemplar memory, converging to the training data becomes difficult, especially for the continual learning model. While a simple solution would be to extend the training epochs for convergence, it would collide with our aim for cost-efficiency. On the other hand, joint-trained models relatively converge better. This result aligns with our conjecture on plasticity (see Section 3.2), the results of [13] that a model sequentially trained on different tasks (i.e., continual learning) suffers a loss of plasticity (i.e., a model’s ability to learn new tasks), as well as the results of [55], which demonstrated that learning a new task becomes difficult under extreme (i.e., full ratio) replay memory settings. We believe a thorough investigation of this new phenomenon is required.
@@ A.5 Experimental Details
In this section, we report the experimental details of our experiments. In all our class-IL experiments, we have used the PyCIL [78] library, which allows easy replication. We followed the standard training hyperparameters of the PyCIL library, which are fixed across experiments. For the LLM continual instruction tuning experiments, we have used the TRACE [64] library. We followed the default training hyperparameters of the TRACE library. Regarding unique hyperparameters, the average interval $j$ was set as $5$, and the warming epoch $n_{warm}$ was set as $25\%$ of the total training epochs (default set as $70$ under the PyCIL setting). $j$ and $n_{warm}$ were selected using a grid search. The hyperparameters searched in the CIFAR-100 benchmark were applied without modification to the ImageNet-100 experiments. In the LLM experiments, the hyperparameters were selected in the $0.2$ ($20\%$) memory setting and applied to the other settings. Note that in the TRACE setting, only one epoch is provided in the replay stage, hence, the average interval was set as $20$ iterations, not epochs. Please refer to Section 4 for a better understanding of each hyperparameter. Regarding the model architectures, we used a ResNet-32 for the CIFAR-100 and a ResNet-18 for the ImageNet-100 experiments, following standard settings in PyCIL. For the TRACE experiments, we used a Llama-3.2-1B model. Lastly, regarding the computing resources, we used a single NVIDIA RTX 6000 GPU for all class-IL experiments. For the LLM experiments, we used three V100 GPUs. For our experiments, we used the 2.2.1 version of Pytorch [52].
@@
Weight Space Operations
Recent works have shown that manipulating model parameters directly with weight-space operations (e.g., model merging [66]) can handle multi-task learning [72] and continual learning [46, 45] in a more principled way. These techniques usually intervene post-training by merging the weights of different models e.g., [68] suggested a selective merging algorithm that mitigates the interference between different models, while [27] showed that arithmetic operations on the weight space can edit the model without training. Unlike these post-training interventions, our approach manipulates the model’s weight space amidst training [28, 29] without storing multiple model parameters, aiming for cost-effective editing of the continual learner. Another relevant yet overlooked topic is the effect of weight-space operations on model attributes e.g., loss landscape [40, 32] and plasticity [13], that contribute to continual learning and generalization. This work empirically investigates various aspects of the model to study their effect on the model’s ability to handle distribution shifts. In the continual learning literature, several works have adopted weight-space operations to obtain multi-task solutions without joint training. For instance, [37] has suggested the use of weight averaging techniques for continual learning, and [45] has extended the idea using task arithmetic. However, these approaches merge models post-training and require the storage of multiple model weights during training. On the other hand, our approach utilizes weight-space operations amidst training, without the redundant storage of multiple model weights. We view this as an important difference in modern settings where the model size is exponentially growing.
A.7 Societal Impacts

### v5

@@ A.4 Experimental Results and Analysis
Table 7: Average class‑IL accuracy (%) on CIFAR‑100 with varying exemplar‑memory sizes. We report experimental results with varying memory sizes, ranging from 20 exemplars per class (a common setting in class-IL) to 500 exemplars per class (storing the entire previous dataset in CIFAR-100. Bold highlights the best non‑expansion method. We report the standard error across 5 runs.
@@
The results of the CIFAR-100 class-IL experiment are reported in Table˜7. Here, we validate that our method is indeed the strongest among non-expansion methods, and even surpasses expansion-based methods under abundant exemplar memory (see Figure˜7(a)). Specifically, we see that in cases where the exemplar memory size is larger than $16\%$ of the full data, our weight space consolidation method outperforms all non-expansion methods, and begins to match the costly expansion-based methods when the exemplar memory ratio exceeds $40\%$. The true strength of our method lies in its training cost (see Figure˜7(b)), where our method’s train time is at par with the cheapest baseline (Replay), while taking one-fifth the time of the expansion-based FOSTER (Wang et al., 2022) and non-expansion-based BiC (Wu et al., 2019).
@@
Table 8: Average class‑IL accuracy (%) on ImageNet‑100 with varying exemplar‑memory sizes. Bold highlights the best non‑expansion method. We report the standard error across 5 runs.
@@
The experimental results in the ImageNet-100 benchmark display a similar pattern. In Figure˜8(a), we observe a gradual increase in average task accuracy as the exemplar memory size increases, which eventually saturates as it enters the abundant exemplar memory regime. The cost of training increases proportionally to the memory size, where methods like BiC or DER require substantially larger training time. On the other hand, our method displays high accuracy while using roughly one-third   half of the training time (see Figure˜8(b)). For detailed results, please refer to Table˜8.
@@
TRACE (Wang et al., 2023) is a continual instruction tuning benchmark for the evaluation of LLMs across eight sequential domains, including science (Lu et al., 2022), policy (Shah et al., 2023), meeting summarization (Hu et al., 2023), multilingual classification (Zhao et al., 2023; Gonzales et al., 2021), code generation (Lu et al., 2021), and math reasoning (Mishra et al., 2022).
@@
Table˜9 reports the LLM continual instruction tuning benchmark results in TRACE. Similar to the results in the Class-IL benchmarks (e.g., CIFAR-100, ImageNet-100), we observe a similar pattern where the accuracy grows in accordance with the increase in memory size (see Figure˜10). An interesting observation is that offline merging (i.e., model merging methods that merge post-training) tends to underperform, displaying lower scores while requiring the VRAM cost of storing multiple models. This aligns with the empirical results of Dziadzio et al. (Dziadzio et al., 2024), which showed that offline merging methods face difficulties in accumulating the multi-task knowledge. Compared to this, our method In addition, note that for TRACE, we use a different method to measure the train cost, which is the relative VRAM usage. The relative VRAM usage is a scaled version of the training time, which measures the computation cost of the training based on the VRAM usage. This scaling is required to distinguish from models that take a similar time to train, but use different numbers of GPUs. For instance, our method and MagMax require similar time to train, but MagMax requires multiple ($t$) task vectors to be loaded in the VRAM, and this is reflected in the relative VRAM usage measure. Considering this, we visualize the results in Figure˜3.
@@
Table 9: General language understanding and reasoning scores on TRACE with varying exemplar-memory sizes. We report the standard error across 5 runs.
@@
Method # Mem. Size # of $\theta$ stored Runtime (min.) C-STANCE FOMC MeetingBank Py150 ScienceQA $\text{NumGLUE}_{\text{cm}}$ $\text{NumGLUE}_{\text{ds}}$ 20Minuten Avg. Acc. Acc. ROUGE-L Similarity Acc. Acc. Acc. SAR I Fine-tune 137’5 0.42 0.25 0.33 0.49 0.21 0.28 0.51 0.37 0.35 Replay 0.2 314’7 0.43 ±0.08 0.01±0.05 0.35±0.08 0.50±0.07 0.29±0.07 0.32±0.06 0.38±0.05 0.37±0.06 0.33 0.4 351’2 0.49±0.09 0.21±0.07 0.38±0.07 0.50±0.05 0.47±0.04 0.35±0.05 0.41±0.06 0.39±0.05 0.40 0.6 394’7 0.47±0.06 0.13±0.05 0.38±0.05 0.55±0.03 0.18±0.04 0.40±0.03 0.56±0.03 0.37±0.04 0.38 0.8 449’4 0.49±0.03 0.13±0.05 0.37±0.03 0.57±0.04 0.21±0.03 0.38±0.03 0.56±0.03 0.37±0.03 0.38 SLERP 0.2 1 375’9 0.42±0.07 0.45±0.05 0.31±0.06 0.54±0.07 0.36±0.03 0.29±0.05 0.42±0.04 0.40±0.04 0.39 0.4 1 422’4 0.45±0.06 0.42±0.06 0.26±0.03 0.53±0.06 0.33±0.05 0.29±0.05 0.48±0.06 0.43±0.06 0.40 0.6 1 472’9 0.42±0.06 0.56±0.05 0.38±0.05 0.57±0.03 0.30±0.03 0.34±0.04 0.54±0.05 0.38±0.06 0.43 0.8 1 526’4 0.45±0.05 0.18±0.05 0.37±0.06 0.58±0.06 0.39±0.04 0.37±0.03 0.53±0.04 0.38±0.04 0.41 Avg. (Online) 0.2 1 374’5 0.44±0.05 0.46±0.06 0.33±0.04 0.52±0.05 0.39±0.06 0.31±0.05 0.37±0.05 0.37±0.06 0.40 0.4 1 421’7 0.42±0.05 0.41±0.05 0.35±0.05 0.55±0.03 0.43±0.07 0.28±0.03 0.43±0.03 0.44±0.05 0.41 0.6 1 471’4 0.45±0.05 0.50±0.05 0.41±0.04 0.55±0.04 0.35±0.05 0.34±0.03 0.52±0.04 0.39±0.05 0.44 0.8 1 524’5 0.45±0.03 0.49±0.05 0.35±0.03 0.52±0.03 0.16±0.04 0.35±0.03 0.47±0.05 0.36±0.04 0.39 Avg. (Offline) 0.2 $t$ 377’1 0.48±0.03 0.02±0.01 0.37±0.03 0.58±0.02 0.29±0.03 0.34±0.05 0.51±0.05 0.37±0.03 0.36 0.4 $t$ 430’6 0.45±0.04 0.03±0.01 0.38±0.02 0.58±0.05 0.33±0.03 0.40±0.03 0.48±0.03 0.37±0.04 0.38 0.6 $t$ 479’3 0.44±0.04 0.56±0.03 0.38±0.04 0.59±0.02 0.36±0.03 0.38±0.04 0.53±0.03 0.36±0.01 0.45 0.8 $t$ 525’4 0.50±0.04 0.21±0.03 0.38±0.03 0.59±0.03 0.35±0.03 0.38±0.02 0.49±0.04 0.37±0.02 0.41 Task-Arith. 0.2 $t$ 374’6 0.51±0.03 0.02±0.01 0.39±0.05 0.56±0.03 0.29±0.04 0.40±0.03 0.52±0.04 0.38±0.05 0.38 0.4 $t$ 427’6 0.49±0.03 0.01±0.01 0.39±0.05 0.59±0.04 0.32±0.05 0.40±0.02 0.53±0.04 0.39±0.03 0.39 0.6 $t$ 474’9 0.50±0.03 0.47±0.06 0.38±0.05 0.60±0.03 0.32±0.03 0.42±0.03 0.52±0.03 0.40±0.03 0.45 0.8 $t$ 523’1 0.49±0.03 0.28±0.03 0.39±0.03 0.57±0.04 0.40±0.03 0.44±0.05 0.48±0.02 0.38±0.03 0.43 MagMax 0.2 $t$ 373’3 0.49±0.04 0.44±0.03 0.25±0.06 0.26±0.06 0.37±0.04 0.31±0.03 0.48±0.04 0.39±0.04 0.37 0.4 $t$ 424’8 0.46±0.04 0.41±0.03 0.39±0.05 0.26±0.06 0.35±0.06 0.43±0.04 0.43±0.05 0.40±0.05 0.39 0.6 $t$ 469’9 0.59±0.05 0.45±0.04 0.37±0.04 0.32±0.06 0.43±0.05 0.42±0.04 0.45±0.03 0.40±0.03 0.42 0.8 $t$ 525’6 0.32±0.03 0.54±0.03 0.23±0.04 0.25±0.06 0.32±0.03 0.29±0.04 0.51±0.03 0.39±0.04 0.36 Ours 0.2 1 370’9 0.39±0.05 0.13±0.03 0.36±0.07 0.51±0.07 0.32±0.04 0.40±0.02 0.56±0.03 0.39±0.01 0.38 0.4 1 422’6 0.41±0.03 0.48±0.02 0.38±0.05 0.55±0.05 0.45±0.02 0.41±0.03 0.30±0.03 0.39±0.02 0.42 0.6 1 465’8 0.44±0.02 0.57±0.02 0.35±0.04 0.52±0.05 0.47±0.02 0.38±0.02 0.55±0.03 0.39±0.01 0.46 0.8 1 522’9 0.46±0.02 0.47±0.02 0.37±0.05 0.55±0.03 0.46±0.02 0.36±0.02 0.53±0.01 0.40±0.01 0.45
Zero-Memory Setting
[…]
Table 10: Average class-IL accuracy (%) on CIFAR-100 under zero memory.
Memory Size Method 0 DER $53.37$ FOSTER $26.43$ MEMO $43.98$ Replay $26.07$ iCaRL $29.03$ BiC $28.29$ WA $42.10$ Ours $29.83$
Sparsity analysis after reset.
Next, we examine how much of the model contributes to learning new tasks. In Figure 11, we vary the retain rate $Q$ in the reset step. We find that resetting up to 80% of parameters yields minimal degradation in accuracy, suggesting that only a small subset of weights are actively involved in continual learning. This aligns with findings in sparse training and pruning (Frankle and Carbin, 2019; Chen et al., 2020). We observe that weight averaging helps retain performance under extreme resets (e.g., $Q{=}10\%$), suggesting that it encourages robust, sparse models that do not rely on a small set of critical parameters.
Effect of Sampling techniques
Table 11: Average class-IL accuracy (%) on CIFAR-100 with varying exemplar-memory sizes using reservoir sampling. Memory Size indicates the number of exemplars per class. The value inside the parentheses indicates the gap between the default random sampling and the reservoir sampling results. Bold highlights the best non‑expansion method.
$\text{Memory Size}_{(\text{ratio of memory to full data})}$
Method $20_{(4\%)}$ $200_{(40\%)}$
DER $64.21_{(+0.26)}$ $74.83_{(+0.19)}$
FOSTER $65.97_{(-0.25)}$ $73.71_{(+0.18)}$
MEMO $62.36_{(+0.37)}$ $73.40_{(-0.31)}$
Replay $49.32_{(+0.69)}$ $71.66_{(+0.06)}$
iCaRL $50.34_{(+0.39)}$ $72.89_{(+0.20)}$
BiC $53.86_{(+0.21)}$ $69.27_{(+0.12)}$
WA $\textbf{61.68}_{(+0.36)}$ $71.65_{(+0.23)}$
Ours $52.49_{(+0.33)}$ $\textbf{74.63}_{(+0.14)}$
In this section, we study the effect of batch sampling in exemplar-based continual learning. Prior work has examined the role of exemplar selection and sampling, with early results suggesting that sophisticated schemes such as reservoir sampling may improve performance under severely constrained memory Kim et al. (2020), but more recent large-scale studies report that random and advanced sampling strategies behave similarly in practice Masana et al. (2022). We expect this trend to persist in the abundant-memory regime considered in our work. To verify this, we re-run all methods on CIFAR-100 using reservoir sampling for exemplar selection, ensuring that the memory buffer fairly represents past tasks. As shown in Table 11, the differences with respect to our default random sampling (values in parentheses) are consistently small, and the gap shrinks further as memory size (i.e., number of exemplars per class in slass-IL) increases from $4\%$ to $40\%$. This indicates that, when sufficient memory is available, the choice of sampling strategy has only a marginal impact on overall performance.
An alternative way to improve plasticity is to enforce a fixed ratio of current-task data within each mini-batch by concatenating current and replay examples at the batch level. While such a design can indeed enhance plasticity, it directly reduces stability: aggressively prioritizing current data accelerates catastrophic forgetting of previous tasks and increases GPU cost due to more constrained batch construction. Thus, batch-level concatenation is complementary but cannot replace our approach, which targets the plasticity–stability trade-off through weight-space operations rather than through carefully engineered batches. Moreover, as highlighted by Dohare et al. (2024), loss of plasticity is a general phenomenon in deep continual learning; our Weight Space Consolidation provides a robust way to restore plasticity that is largely independent of specific batch-design or sampling heuristics.
Effect of Longer Task Sequences
Table 12: Average class-IL accuracy (%) on CIFAR-100 with varying exemplar-memory sizes on longer task sequences (T=20). Memory Size indicates the number of exemplars per class. The value inside the parentheses indicates the gap between the default task sequence (T=10; 10 tasks) and the longer task sequence setting (T=20; 20 tasks).
$\text{Memory Size}_{(\text{ratio of memory to full data})}$
Method $20_{(4\%)}$ $200_{(40\%)}$
DER $55.42_{(-8.53)}$ $70.91_{(-3.73)}$
FOSTER $55.94_{(-10.28)}$ $71.80_{(-1.73)}$
MEMO $54.61_{(-7.38)}$ $70.14_{(-3.57)}$
Replay $44.73_{(-3.90)}$ $70.88_{(-0.72)}$
iCaRL $45.28_{(-4.67)}$ $69.24_{(-3.45)}$
BiC $45.15_{(-8.50)}$ $68.97_{(-0.18)}$
WA $\textbf{52.36}_{(-8.96)}$ $70.35_{(-1.07)}$
Ours $49.77_{(\textbf{-2.39})}$ $\textbf{76.65}_{(\textbf{+0.22})}$
To assess how our findings scale with the length of the continual learning stream, we further evaluate all methods on a longer task sequence with $T=20$ tasks on CIFAR-100, while keeping the total data and exemplar-memory budget fixed. As shown in Table 12, average class-IL accuracy systematically degrades when moving from the default $T=10$ setting to $T=20$, and this degradation is particularly pronounced in the low-memory regime ($20$ exemplars per class), indicating that longer task sequences amplify both forgetting and the loss of plasticity. Nonetheless, in the sufficient-memory setting ($200$ exemplars per class, i.e., $40\%$ memory), our method maintains strong performance and exhibits a comparatively smaller drop than many baselines, demonstrating that our Weight Space Consolidation remains effective even as the task sequence becomes longer.
@@
Lastly, we show that under abundant exemplar memory, convergence becomes difficult, similar to a multi-task learning setting. The results are illustrated in Figure˜2(b). Specifically, we analyze the training loss of the model under two cases. (1) Continual Learning: train the model across 20 tasks sequentially, with abundant exemplar memory. (2) Full retraining (Joint Training): train individual models for each task using all known task data. For this experiment, we used the CIFAR-100 dataset divided into 20 subtasks. In Figure˜2(b), we can see that under abundant exemplar memory, converging to the training data becomes difficult, especially for the continual learning model. While a simple solution would be to extend the training epochs for convergence, it would collide with our aim for cost-efficiency. On the other hand, joint-trained models relatively converge better. This result aligns with our conjecture on plasticity (see Section˜3.2), the results of (Dohare et al., 2024) that a model sequentially trained on different tasks (i.e., continual learning) suffers a loss of plasticity (i.e., a model’s ability to learn new tasks), as well as the results of (Rolnick et al., 2019), which demonstrated that learning a new task becomes difficult under extreme (i.e., full ratio) replay memory settings. We believe a thorough investigation of this new phenomenon is required.
@@ A.5 Experimental Details
In this section, we report the experimental details of our experiments. In all our class-IL experiments, we have used the PyCIL (Zhou et al., 2023a) library, which allows easy replication. We followed the standard training hyperparameters of the PyCIL library, which are fixed across experiments. For the LLM continual instruction tuning experiments, we have used the TRACE (Wang et al., 2023) library. We followed the default training hyperparameters of the TRACE library. Regarding unique hyperparameters, the average interval $j$ was set as $5$, and the warming epoch $n_{warm}$ was set as $25\%$ of the total training epochs (default set as $70$ under the PyCIL setting). $j$ and $n_{warm}$ were selected using a grid search. The hyperparameters searched in the CIFAR-100 benchmark were applied without modification to the ImageNet-100 experiments. In the LLM experiments, the hyperparameters were selected in the $0.2$ ($20\%$) memory setting and applied to the other settings. Note that in the TRACE setting, only one epoch is provided in the replay stage, hence, the average interval was set as $20$ iterations, not epochs. Please refer to Section˜4 for a better understanding of each hyperparameter. Regarding the model architectures, we used a ResNet-32 for the CIFAR-100 and a ResNet-18 for the ImageNet-100 experiments, following standard settings in PyCIL. For the TRACE experiments, we used a Llama-3.2-1B model. Lastly, regarding the computing resources, we used a single NVIDIA RTX 6000 GPU for all class-IL experiments. For the LLM experiments, we used three V100 GPUs. For our experiments, we used the 2.2.1 version of Pytorch (Paszke et al., 2019).
@@
Weight Space Operations.
Recent works have shown that manipulating model parameters directly with weight-space operations (e.g., model merging (Wortsman et al., 2022)) can handle multi-task learning (Yu et al., 2024) and continual learning (Marouf et al., 2025; Marczak et al., 2025) in a more principled way. These techniques usually intervene post-training by merging the weights of different models e.g., (Yadav et al., 2024) suggested a selective merging algorithm that mitigates the interference between different models, while (Ilharco et al., 2022) showed that arithmetic operations on the weight space can edit the model without training. Unlike these post-training interventions, our approach manipulates the model’s weight space amidst training (Izmailov et al., 2018; Jang et al., 2025) without storing multiple model parameters, aiming for cost-effective editing of the continual learner. Another relevant yet overlooked topic is the effect of weight-space operations on model attributes e.g., loss landscape (Li et al., 2018; Kaur et al., 2023) and plasticity (Dohare et al., 2024), that contribute to continual learning and generalization. This work empirically investigates various aspects of the model to study their effect on the model’s ability to handle distribution shifts. In the continual learning literature, several works have adopted weight-space operations to obtain multi-task solutions without full retraining. For instance, (Kozal et al., 2024) has suggested the use of weight averaging techniques for continual learning, and (Marczak et al., 2025) has extended the idea using task arithmetic. However, these approaches merge models post-training and require the storage of multiple model weights during training. On the other hand, our approach utilizes weight-space operations amidst training, without the redundant storage of multiple model weights. We view this as an important difference in modern settings where the model size is exponentially growing.
A.7 Limitations
While our method demonstrates strong performance and cost-efficiency under abundant exemplar memory, it assumes access to a representative subset of past data, which may not always be feasible in privacy-sensitive or streaming-only environments. Additionally, our analysis primarily focuses on class-incremental learning and continual instruction tuning with relatively clean task boundaries. Future work may explore how the proposed weight-space strategies generalize to more complex settings such as task-agnostic CL, online CL, or continual reinforcement learning.
A.8 Societal Impacts